In a recent survey, a researcher claimed that the average life expectancy of people in a certain country is 72 years. Is his claim correct if a random sample of 25 deaths from his country showed a mean of 71.2 years with standard deviation of 1.2 years? Use 99% confidence level.



1
Expert's answer
2022-06-07T09:22:30-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=72"

"H_1:\mu\not=72"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=24" and the critical value for a two-tailed test is "t_c =2.79694."

The rejection region for this two-tailed test is "R = \{t:|t|>2.79694\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{71.2-72}{1.2/\sqrt{25}}=-3.33333"


Since it is observed that "|t|=3.33333>2.79694=t_c," it i s then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=24" degrees of freedom, "t=-3.33333" is "p=0.002776," and since "p= 0.002776<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 72, at the "\alpha = 0.01" significance level.